A cook uses a microwave oven to heat a meal. The radiation wavelength is 9.25 cm. What is the energy of one photon of this microwave radiation?


1
Expert's answer
2022-02-15T09:58:13-0500

Given:

"\lambda=\rm 9.25\: cm=9.25*10^{-2}\: m"


The energy of one photon of this microwave radiation

"E=\frac{hc}{\lambda}\\
=\rm \frac{6.63*10^{-34}\: J\cdot s*3*10^8\: m/s}{9.25*10^{-2}\: m}=2.15*10^{-24}\: J"